OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1

OpenStudy (anonymous):

First, solve for x in the first equation. 2x + y = 1 so, we get 2x = -y + 1 and then x = -y/2 + 1/2

OpenStudy (anonymous):

There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

2x + y = 1 4x + 2y = −1 rearrange the first like this: y = 1 -2x now sub in this as y in the second: 4x + 2(1-2x) = -1 4x - 4x + 2 = -1 2 = -1 hence inconsistent. no solutions

OpenStudy (anonymous):

Here are my choices: x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

So it would be C?

OpenStudy (anonymous):

2x + y = 1 is the same as y = 1 - 2x 4x + 2y = −1 is the same as y = -0.5 - 2x

OpenStudy (anonymous):

So D?